OpenStudy (anonymous):

6t^2-5.4-3.5t+2.5t^2+3t

OpenStudy (anonymous):

8.5t^2-0.5t-5.4

OpenStudy (anonymous):

Let's simplify step-by-step. 6t2−5.4−3.5t+2.5t2+3t =6t2+−5.4+−3.5t+2.5t2+3t Combine Like Terms: =6t2+−5.4+−3.5t+2.5t2+3t =(6t2+2.5t2)+(−3.5t+3t)+(−5.4) =8.5t2+−0.5t+−5.4 Answer: =8.5t2−0.5t−5.4